OpenStudy (anonymous):

How much work is done on an object that travels 3 meters under a constant force of 2 Newtons?

OpenStudy (anonymous):

the formula for work is W = F.d (force * displacement) SO just substitute , and u get W = 2*3 = 6J

OpenStudy (anonymous):

Work = force * distance Work = 2N * 3m = 6 N*m = 6.0 Joules of work done